OpenStudy (aivantettet26):

A 1370-kg car is skidding to a stop along a horizontal surface. The car decelerates from 27.6 m/s to a rest position in 3.15 seconds. Assuming negligible air resistance, determine the coefficient of friction between the car tires and the road surface.

OpenStudy (anonymous):

[ F_f = \mu F_n\] Normal Force \[ F_n = mg = 1370 * 9.81 = 13439.7 N\] Friction Force \[ F_f = ma\] car decelerates from 27.6 m/s to 0 in 3.15 sec \[ a = \frac{dv}{dt} = \frac{27.6-0}{3.15-0} = 8.7619 \space \frac{m}{s^2} \] therefor: \[ F_f = 1370 * 8.7619 = 12003.8 N\] and \[ \mu = \frac{F_n}{F_f} = \frac{13439.7}{12003.8} = 1.1196\]

OpenStudy (aivantettet26):

the answer is 0.894

OpenStudy (fifciol):

\[F_{fr}=ma\] \[\mu mg=ma\] \[\mu g=\frac{ v }{ t }\] \[\mu=\frac{ v }{g t }=0,894\]

OpenStudy (anonymous):

the last part i did is wrong: u = Ff/Fn = 12003.8/134349.7 = 0.8931

OpenStudy (fifciol):

yes , friction force is proportional to normal force

OpenStudy (aivantettet26):

thank you all!